# A mathematical theory of the critical point of ferromagnetic Ising systems

### v1

@@ 1.3.1 Generalities
Equation (1.3.3) is due to the stability condition (1.2.4). We define the spin-flip symmetry $\alpha$ by the following equation, valid for all $x\in\mathbb{Z}^{\nu}$:
@@
The case of zero anisotropy corresponds to the Ising model. We shall consider this boundary case in most of the paper, but follow ([BRo297], p.320 et seq.) in considering the associated states as states on the whole algebra $\mathcal{A}$ as before. This choice agrees with the physical significance of the model in applications, which relies on the fact that it is an anisotropic limit ($\delta\rightarrow 0$) of the model (1.2.1). The dynamics of Subsection 3.2 also takes this choice into account. When restricting to observables which are functionals of the $\sigma{}_{x}^{3}$; $x\in\mathbb{Z}^{\nu}$, however, the correlation inequalities which are specific to the Ising model, remain, of course, applicable, as well as the probabilistic language, commonly associated with the “classical version”.
Theorem 1.2 substantiates considering the model defined by equations (1.2.1) as the Ising universality class. An additional reason for this nomenclature is the property of Ising domination proved in Appendix A.
Although we keep the dimension $\nu$ general, for clarity, we shall only consider , when dealing with short-range interactions, the (physically relevant) cases $\nu=2$ and $\nu=3$ in correspondence with Theorem 1.2. The critical dimension is $\nu_{c}=4$, as recently shown in [ADC21], and for $\nu\geq\nu_{c}$ it is believed that there are no anomalous fluctuations (the exponent $\eta=0$), and that mean-field (Gaussian) behavior takes place.
@@ 1.3.2 A criterion of criticality: Criterion A
The state $\bar{\omega}_{T}$ is spin-flip and translation invariant, and, by Theorem  1.2 it follows immediately that
@@ 1.3.5 Interacting and quasi-free theories
The thermal Wightman functions of the Ising model ((1.2.1) with $\delta=0$) are expected to define an interacting system in the sense of Definition  1.3, except if $T=0$. In this case, the ergodic states are (infinite) product states, see also Appendix A. Otherwise, for $T$ sufficiently low (nonzero) the connected four-point Wightman function is presumably not identically zero: it is expected that this fact may be proved by polymer expansions (see [BLOC98] and references given there, in particular their Lemma 1). See also the forthcoming remark.
@@
Regarding the conjecture that the thermal Wightman functions define an interacting system, if we regard Newman’s “free energy” (B.0.2) (Appendix B) as a functional of the state (see the introduction and [Sew86], 3.4.2, p. 60), it is a functional of the equilibrium state, which is unique for $T=T_{c}$ and $h=0$ (here we amplify the discussion to include the external field). Thus, the conjecture is equivalent to the assumption that the Non-Gaussian behavior found in [Ne79] (see also Corollary B.4 of Appendix B) should be true in the Ising case for both $\nu=2$ and $\nu=3$. The condition $p\neq 2$ in Corollary B.4 means that the critical exponent $\delta$ satisfies the condition $\delta\neq 1$. This is true for the Ising model at dimension $\nu=2$ by the rigorous result of Camia, Garban and Newman [CGN14], who showed that $\delta=15$ ($p=1+1/\delta=1+1/15$). For the Ising model in dimension $\nu=3$, the Buckingham-Gunton inequality of Appendix B (equation (B.0.13) implies, together with the forthcoming Glimm-Jaffe inequality (2.3.26), that the condition $\delta\neq 1$ is true whenever the critical exponent $\nu<\infty$. The value $\delta=15$ for the Ising model in dimension $\nu=2$ is given in Table 16.2, p.398, of K. Huang’s book [KH87], while for Ising ($\nu=3$) the value $\delta=5$ is provided, together with $\nu=0.64$. From the same table, experimental values of $\delta$ lie between $4-5$, and those of $\nu$, between $0.6$ and $0.7$.
@@ 1.4 Macroscopic phenomena
represents a classical mixture. This remark applies to both the magnetic and the lattice fluid pictures, of course, but, in the latter case, when the temperature $T\nearrow T_{c}$, the resulting mixture displays an outstanding classical (macroscopic) effect - the critical opalescence - discussed in Subsection 1.1. In the former case, it is known that domains of a magnet are, in nature, found only in a definite direction. This constitutes one of the rare examples observed in nature of the so-called superselection sectors [Wight95] originating from interactions with the environment: a rigorous model of this situation may be constructed precisely around (1.2.1) following the model suggested by Narnhofer and Thirring [NT99], see [Wre23], Section 4.2, and references given there, as well as [NT99] and [NW14]. An alternative approach to this problem was suggested in [MR02], in which it was shown that every generic many-body system contains within the class of microscopic quantum observables a subalgebra of macro-observables, the spectrum of which comprises the macroscopic properties of the many-body system (see also [GL61]). A third, independent, approach to show the existence of domain walls was initiated in [ASW75] and further considerably extended by Koma, Nachtergaele, Spitzer and Starr ([KN97], [KN98], [KNS01], [NSS07]), see also Remark A.1 in Appendix A.
@@
[…]
@@
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{\pm,T}^{\mathrm{e}}(\left.\sigma_{x}^{i}\right/|\Lambda|)=0\qquad\text{ for \quad}i=1,2\ ,$ (1.4.10)
@@
([CW79], III.1.3), where $r^{3}$ denotes the three-dimensional abelian Lie algebra, isomorphic to the Lie algebra of $\mathbb{R}^{3}$. If, in (1.4.8) the factors $1/\sqrt{\left|\Lambda\right|}$ are both replaced with $1/\left|\Lambda\right|^{\alpha}$, with $\alpha>1/2$, i.e., the fluctuations of $\sigma^{1,2}$ are also anomalous, (1.4.12) is also true a fortiori: see [Sew86], Appendix B, p.166, for a general theorem in this connection. Equations (1.4.11) and (1.4.12) together imply that the critical fluctuation state $\omega_{T}^{F}$ of the ferromagnets in the Ising universality class (1.2.1) is characterized, from the point of view of the algebra of fluctuations, as the unique state for which the Lie algebra of fluctuations is classical.
@@
A final, most important, aspect of the relation between the states $\omega_{T}^{F}$ and $\omega_{T}$ is the possible occurrence of a basic structural difference between them, taking into account, of course, that they are states on different algebras of observables, $\mathcal{P}$ and $\mathcal{A}$, as previously remarked. One possibility, proved under certain natural assumptions in the forthcoming Theorem  2.6, is that $\omega_{T}$ is an interacting state, while $\omega_{T}^{F}$ is a quasi-free state. We shall now try to provide some intuition, why this may be expected to occur.
The bosonic quasi-spin operators (1.4.5), (1.4.6) (spin-waves, with $\left|m_{\pm}\right|=1$) have been studied in great detail by Wehrl in the BCS model [We71]. In a certain limit (the Trotter limit, also used in [CW99], [CW79]), their dynamics is shown to satisfy the KMS condition (1.2.13) only if the value $k=0$ is excluded in (1.4.5), (1.4.6). The zero-momentum fluctuation operators are invariant under all permutations of the elements of the set $\{x\in\Lambda\}$, as is his Hamiltonian (which is the BCS mean-field Hamiltonian), and, comparing this situation with ordinary many-body systems, one sees that the operators $a_{\Lambda}$, $a_{\Lambda}^{\ast}$, in (1.4.2), (1.4.3), are the analogues of the center of mass variables, and thus behave freely. Consequently, they do not “thermalize”, and this happens for all temperatures. If we believe that these considerations apply, qualitatively speaking, to general models, we are led to conjecture that, for all temperatures (including $T_{c}$), the fluctuation state is quasi-free.
@@ 2.2 The case $0<T<T_{c}$
Theorem 3.3 of [MR02] may now be stated; we briefly sketch the proof for the reader’s convenience, because the arguments will be of major importance when dealing with the critical point:
@@ 2.3 The case of anomalous fluctuations: the critical (ergodic) state for $r=2$ and $T=T_{c}$
where the superscript ${}^{\vee}$ denotes the inverse of the Fourier transform. We have, however, replaced their $\alpha$ by our $\rho$, which satisfies (2.3.16), i.e., their assumption. Our conventions for the Fourier transform differ from those in [LL00], so that the above formulas may differ from those in our notation by constants which do not alter the final statement. If we now set
@@
It is the necessity of using a definite normalization for all $r$ in (2.4.4) – the only one which guarantees the existence of the thermodynamic limit for $r=2$, due to the existence of anomalous fluctuations – which leads to a quasi-free fluctuation state in the above theorem; in particular, the fact that the case of equality (2.4.11) satisfies (2.4.5).
@@ 3 Long-range interactions: universality and its violations
3.1 Universality and the problem with s.s.b. of continuous symmetries
It is well-known that the critical exponents or indices depend only on rather general qualitative features of physical systems, such as dimensionality, symmetry and range of interaction: this is the universality of critical phenomena [Fis67]. We shall be concerned with the two issues of symmetry and interaction range, and consider the case of s.s.b. of a continuous symmetry (the rotation symmetry). One important example is the isotropic Heisenberg antiferromagnet ((1.2.1), with $j(|x|)\leq 0$, a nearest-neighbor interaction, $h=0$ and $\delta=1$, for which a phase transition has been proved [FILS78], with s.s.b. [KT93]. It turns out, however, that, for this model and others, covered by the results of [LPW81], Criterion A does not apply, due to the Goldstone phenomenon (see also [MR02]):
Proposition 3.1.
In the case of s.s.b. of a continuous symmetry, Criterion A does not apply.
Proof. By Theorem 1 (b) of [LPW81], ergodic states for $T<T_{c}$ cannot exhibit $l^{1}$- clustering, and, thus, the critical state is not singled out by Criterion A.
$\Box$
For $T=0$ the analogue of Proposition 3.1 is the Goldstone theorem: it states that the spectrum of the physical Hamiltonian is gapless [WFW87]. This situation corresponds to the existence of zero mass particles in quantum field theory. In the present model the analogous fact is that the inverse correlation length $c_{T}$ in (1.3.6) is expected to be zero. But this is expected to happen  only at $T=T_{c}$, and not throughout the whole transition region $0\leq T\leq T_{c}$. This is a basic open problem (see also the conclusion).
The result of [LPW81] was generalized by Martin [Mar82] in an optimal way to include several types of decay of the interaction potential, in particular the case of dimension $\nu=2$, providing an extension of the Hohenberg-Mermin-Wagner theorem ([H67], [MeW66]). The latter theorem precludes a first-order phase transition (i.e., with nonzero spontaneous magnetization) in n-vector models, with $SO(n)$ symmetry, but such have been rigorously proved to exist by van Enter and Shlosman [vES02] with sufficiently nonlinear interaction potentials. The importance of this result pertains to the general subject of the limitations of the universality concept just stated, in that it provides explicit examples which show that classes of interactions exist, all elements of which have the same kind of phase transition between a high temperature and a low-temperature phase, which are not determined only by the dimension of the system, the symmetry of the interaction and the long-range or short-range character of the interaction: indeed, in their case, the degree of nonlinearity also plays a decisive role.
On the other hand, given a dimension and a symmetry class, the short vs. long range character of the interaction may be of decisive importance in determining whether a phase transition of a given type (e.g., first-order) takes place in a system, e.g. for Ising systems in dimension $\nu=1$ no first-order phase transition is possible if the interaction is short-range; in order to have one, we must consider long-range interactions, to which we now turn.
3.2 Long-range interactions
3.2.1 Brief review of results and problems
@@
This is not quite (3.2.3), but it is believed that the result is correct, in agreement with the predictions of Bleher and Sinai ([BleSi73], [BleSi75], see also [Sew86], Sec. 5.4.3 for a pedagogic outline) for the Dyson hierarchical model. We come back in Section 4 to a more detailed analysis of their predictions.
@@
3.2.2 Aspects of dynamics: violations of universality in problems of energetic and dynamic stability
@@
In the critical region (3.2.3), this energy is of an order of magnitude greater than that which is necessary for short range interactions (e.g., bounded by a constant in the case of finite range). This is an expression of Haag’s principle for the Dyson model, whose validity is, however, restricted to the values of $\alpha$ corresponding to the phase-transition region. This is not an accident: (3.2.7) is the central ingredient in an energy-entropy estimate which may even be used to prove the existence of a phase transition in the model (see [Sew86], p. 126 and references given there). If one considers now a state consisting of a large droplet of spins polarized along the $1$-direction, immersed in a background of spins polarized along the $3$-direction, the energy fluctuation $\Delta E$ of the Dyson Hamiltonian in this state is of the order of the r.h.s. of (3.2.7). This state will be unstable and decay with a half-life $\Gamma$ given by the time-energy uncertainty principle $\Delta E\cdot\Gamma\approx\hbar$ (in the case of a Markovian approach to equilibrium , i.e., of type $\exp(-t/\Gamma)$), otherwise, instead of $\Gamma$, one would have a quantity called “sojourn time” in the state (see, e.g., [MWB13], Theorem 3.17, p. 81, and references given there). From this heuristics and (3.2.7), we are led to conclude: the rate of decay of such a state is the greater, the longer the range of interaction, i.e, the closer the parameter $\alpha$ is to one, inside the open interval $(1,2)$. We shall call this heuristic principle, for brevity, “Haag’s principle”, because it is a direct consequence of the rise of energetic stability of large droplets when the range of interaction increases. If, in (1.2.1), we consider the Ising limit, i.e., set $\delta\equiv 0$, we obtain the generalized Ising model (gIm) studied by Emch [Em66] and Radin [Ra70] (we changed the overall sign to agree with Radin’s notation)

### v2

@@ 1.3.1 Generalities
Equation (1.3.3) is due to the stability condition (1.2.4). We define the spin-reflection symmetry $\alpha$ by the following equation, valid for all $x\in\mathbb{Z}^{\nu}$:
@@
The case of $\delta=0$ corresponds to the Ising model. We shall consider this boundary case in most of the paper, but follow ([BRo297], p.320 et seq.) in considering the associated states as states on the whole algebra $\mathcal{A}$ as before. This choice agrees with the physical significance of the model in applications, which relies on the fact that it is an anisotropic limit ($\delta\rightarrow 0$) of the model (1.2.1). The dynamics of Subsection 3.2 also takes this choice into account. When restricting to observables which are functionals of the $\sigma_{x}^{3}$; $x\in\mathbb{Z}^{\nu}$, however, the correlation inequalities which are specific to the Ising model, remain, of course, applicable, as well as the probabilistic language, commonly associated with the “classical version”.
Theorem 1.2 substantiates considering the model defined by equations (1.2.1) as belonging to the Ising universality class. An additional reason for this nomenclature is the property of Ising domination proved in Appendix A.
Although we keep the dimension $\nu$ general, for clarity, we shall only consider , when dealing with short-range interactions, the (physically relevant) cases $\nu=2$ and $\nu=3$ in correspondence with Theorem 1.2. An upper bound to the upper critical dimension is $\nu_{c}=4$ [ADC21]. For high dimensions $\eta=0$, and the systems rescales to a “massless” Gaussian, which is quasi-free, but has connected correlations decaying as $r^{2-\nu}$. The very recent paper by Duminil-Copin and Panis [DCP24] proves that the critical exponent $\eta=0$ for $\nu$ at least 4, and gives improved rigorous bounds on $\eta$ for $\nu=3$.
@@ 1.3.2 A criterion of criticality: Criterion A
The state $\bar{\omega}_{T}$ is spin-flip and translation invariant, and, by Theorem 1.2 it follows immediately that
@@ 1.3.4 The connected Wightman functions
Shlosman [Shl86] proved the general statement about the signs for higher-order Ursell functions.
@@ 1.3.5 Interacting and quasi-free theories
The thermal Wightman functions of the Ising model ((1.2.1) with $\delta=0$) are expected to define an interacting system in the sense of Definition  1.3, except if $T=0$. In this case, the ergodic states are (infinite) product states, see also Appendix A. Otherwise, the connected four-point Wightman function is not expected to be identically zero. This is because this would entail that all higher connected functions vanish, and hence already for finite $|\Lambda|$ the characteristic function would be a Gaussian, a behavior not expected for the Ising model which may be viewed as a system of interacting “fermions” in the sense of ([BRo297], p. 426).
@@ 1.4 Macroscopic phenomena
represents a classical mixture. This remark applies to both the magnetic and the lattice fluid pictures, of course, but, in the latter case, when the temperature $T\nearrow T_{c}$, the resulting mixture displays an outstanding classical (macroscopic) effect - the critical opalescence - discussed in Subsection 1.1. In the former case, it is known that domains of a magnet are, in nature, found only in a definite direction. This constitutes one of the rare examples for which the existence of so-called superselection sectors [Wight95] originating from interactions with the environment, which are observed in nature, may be proved: a rigorous model of this situation may be constructed precisely around (1.2.1) following the model suggested by Narnhofer and Thirring [NT99], see [Wre23], Section 4.2, and references given there, as well as [NT99] and [NW14]. An alternative approach to this problem was suggested in [MR02], in which it was shown that every generic many-body system contains within the class of microscopic quantum observables a subalgebra of macro-observables, the spectrum of which comprises the macroscopic properties of the many-body system (see also [GL61]). A third, independent, approach to show the existence of domain walls was initiated in [ASW75] and further considerably extended by Koma, Nachtergaele, Spitzer and Starr ([KN97], [KN98], [KNS01], [NSS07]), see also Remark A.1 in Appendix A.
@@
$\omega_{T}^{F}\left(\left(F_{T}^{i}\right)^{n}\right)=\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{T}\left(\left(F_{\Lambda,T}^{i}\right)^{n}\right)\qquad\text{ with \quad}n=1,2,\ldots\quad\text{and}\quad i=1,2\ .$ (1.4.9)
@@
$\lim_{\Lambda\nearrow\mathbb{Z}^{\nu}}\omega_{\pm,T}^{\mathrm{e}}(\left.\sigma_{x}^{i}\right/|\Lambda|)=0\qquad\text{ for \quad}i=1,2\,$ (1.4.10)
@@
([CW79], III.1.3), where $r^{3}$ denotes the three-dimensional abelian Lie algebra, isomorphic to the Lie algebra of $\mathbb{R}^{3}$. If, in (1.4.8) the factors $1/\sqrt{\left|\Lambda\right|}$ are both replaced with $1/\left|\Lambda\right|^{\alpha}$, with $\alpha>1/2$, i.e., the fluctuations of $\sigma^{1,2}$ are also anomalous, (1.4.12) is also true a fortiori: see [Sew86], Appendix B, p.166, for a general theorem in this connection. Equations (1.4.11) and (1.4.12) together imply that the critical fluctuation state $\omega_{T}^{F}$ of the ferromagnets in the Ising universality class (1.2.1) is characterized, from the point of view of the algebra of fluctuations, as the unique state for which the Lie algebra of fluctuations is classical. This explains why critical behavior in quantum models may be described by classical Ising models.
We refer also to the references on the extensive work on the quantum central limit theorem given in the book by Verbeure [Ve11].
@@
A final, most important, aspect of the relation between the states $\omega_{T}^{F}$ and $\omega_{T}$ is the possible occurrence of a basic structural difference between them, taking into account, of course, that they are states on different algebras of observables, $\mathcal{P}$ and $\mathcal{A}$, as previously remarked. One possibility, proved under certain natural assumptions in the forthcoming Theorem 2.6, is that $\omega_{T}$ is an interacting state, while $\omega_{T}^{F}$ is a quasi-free state. We shall now try to provide some intuition, why this may be expected to occur.
The bosonic quasi-spin operators (1.4.5), (1.4.6) (spin-waves, with $\left|m_{\pm}\right|=1$) have been studied in great detail by Wehrl in the BCS model [We71]. In a certain limit (the Trotter limit, also used in [CW99], [CW79]), their dynamics is shown to satisfy the KMS condition (1.2.12) only if the value $k=0$ is excluded in (1.4.5), (1.4.6). The zero-momentum fluctuation operators are invariant under all permutations of the elements of the set $\{x\in\Lambda\}$, as is its Hamiltonian (which is the BCS mean-field Hamiltonian), and, comparing this situation with ordinary many-body systems, one sees that the operators $a_{\Lambda}$, $a_{\Lambda}^{\ast}$, in (1.4.2), (1.4.3), are the analogues of the center of mass variables, and thus behave freely. Consequently, they do not “thermalize”, and this happens for all temperatures. If we believe that these considerations apply, qualitatively speaking, to general models, we are led to conjecture that, for all temperatures (including $T_{c}$), the fluctuation state is quasi-free.
@@ 2.2 The case $0<T<T_{c}$
Theorem 3.3 of [MR02] may now be stated; we briefly sketch the proof for the reader’s convenience, because the argumements will be of major importance when dealing with the critical point:
@@ 2.3 The case of anomalous fluctuations: the critical (ergodic) state for $r=2$ and $T=T_{c}$
where the superscript denotes the inverse of the Fourier transform. We have, however, replaced their $\alpha$ by our $\rho$, which satisfies (2.3.16), i.e., their assumption. Our conventions for the Fourier transform differ from those in [LL00], so that the above formulas may differ from those in our notation by constants which do not alter the final statement. If we now set
@@
It is the necessity of using a definite normalization for all $r$ in (2.4.4) – the only one which guarantees the existence of the thermodynamic limit for $r=2$, due to the existence of anomalous fluctuations – which leads to a quasi-free fluctuation state in the above theorem; in particular, the fact that the case of equality (2.4.11) satifies (2.4.5).
@@ 3 Long-range interactions: universality and its violations
It is well-known that the critical exponents or indices depend only on rather general qualitative features of physical systems, such as dimensionality, symmetry and range of interaction: this is the universality of critical phenomena [Fis67]. We shall be concerned only with the Ising universality class, which concerns only a discrete symmetry (the spin-reflection symmetry), but consider in this section long-range interactions. Indeed, the short vs. long range character of the interaction may be of decisive importance in determining whether a phase transition of a given type (e.g., first-order) takes place in a system, e.g. for Ising systems in dimension $\nu=1$ no first-order phase transition is possible if the interaction is short-range; in order to have one, we must consider long-range interactions, to which we now turn.
3.1 Brief review of results and problems
@@
This is not quite (3.1.3), but it is believed that the result is correct, in agreement with the predictions of Bleher and Sinai ([BleSi73], [BleSi75], see also [Sew86], Sec. 5.4.3 for a pedagogic outline) for the Dyson hierarchical model. It is of interest that Lohmann, Slade and Wallace [LSW17] have proved that the exponent $\eta$ keeps its mean-field value even for $\alpha>3/2$. We come back in Section 4 to a more detailed analysis of the predictions of Bleher and Sinai.
@@
3.2 Aspects of dynamics: violations of universality in problems of energetic and dynamic stability
@@
In the critical region (3.1.3), this energy is of an order of magnitude greater than that which is necessary for short range interactions (e.g., bounded by a constant in the case of finite range). This is an expression of Haag’s principle for the Dyson model, whose validity is, however, restricted to the values of $\alpha$ corresponding to the phase-transition region. This is not an accident: (3.2.1) is the central ingredient in an energy-entropy estimate which may even be used to prove the existence of a phase transition in the model (see [Sew86], p. 126 and references given there). A proof of phase transitions in the Dyson model along these lines was given in [CFMP05].
If one considers now a state consisting of a large droplet of spins polarized along the $1$-direction, immersed in a background of spins polarized along the $3$-direction, the energy fluctuation $\Delta E$ of the Dyson Hamiltonian in this state is of the order of the r.h.s. of (3.2.1). This state will be unstable and decay with a half-life $\Gamma$ given by the time-energy uncertainty principle $\Delta E\cdot\Gamma\approx\hbar$ (in the case of a Markovian approach to equilibrium , i.e., of type $\exp(-t/\Gamma)$), otherwise, instead of $\Gamma$, one would have a quantity called “sojourn time” in the state (see, e.g., [MWB13], Theorem 3.17, p. 81, and references given there). From this heuristics and (3.2.1), we are led to conclude: the rate of decay of such a state is the greater, the longer the range of interaction, i.e, the closer the parameter $\alpha$ is to one, inside the open interval $(1,2)$. We shall call this heuristic principle, for brevity, “Haag’s principle”, because it is a direct consequence of the rise of energetic stability of large droplets when the range of interaction increases. If, in (1.2.1), we consider the Ising limit, i.e., set $\delta\equiv 0$, we obtain the generalized Ising model (gIm) studied by Emch [Em66] and Radin [Ra70] (we changed the overall sign to agree with Radin’s notation)